# 3-D CMZ I: Central Molecular Zone Overview

### v1

@@ 5.3 The Total Dense Gas Mass of the CMZ
The column density map of the CMZ can be integrated to estimate the total dense molecular gas mass of the CMZ. As described in Section 3, we use the fore/background subtracted maps for our main analysis, but include constraints from the maps with the fore/background included in a few places as well. We consider any gas with a measured column density of N(H2) $>10^{22}$ cm-2 to be dense. In the translation of high column density dust seen with Herschel to the inference of dense, molecular gas are several assumptions: 1) high column density gas corresponds to high volume density gas and 2) dust seen with Herschel traces dense, molecular gas. These assumptions are supported by the many studies (e.g. Mills et al., 2018b; Krieger et al., 2017; Ginsburg et al., 2016) which find abundant dense, molecular gas in the CMZ, with intensity peaks that largely correspond with the Herschel data (Mills & Battersby, 2017). We estimate that most of the gas traced with Herschel observations is likely $n>10^{4}\>\mathrm{cm}^{-3}$ and we refer to this as ‘dense, molecular gas’ throughout.
@@ 5.4 A Ridge of Warm Dust
In the second hypothesis, the ridge of warm dust is thought to be due to heating from young, still-forming or just formed massive stars that follow the morphology of this structure. Detailed maps, including some with temperature, from Hankins et al. (2019, 2020) as well as from the GLIMPSE survey (Benjamin et al., 2003) show mid-IR emission from many young, massive star-forming regions in the CMZ. While there are quite a few located in the central region of the warm dust ridge just above SgrA*, otherwise there seems to be no special association of the young, forming massive star population with the morphology of the dust ridge or other warm spots. In Figure 5 we show the warm dust contours on top of the Spitzer 8 $\mu$m data from the GLIMPSE survey (Benjamin et al., 2003). The central bright massive star-forming regions are clearly seen at 8 $\mu$m, but otherwise, we find no trend with the morphology of the warm dust contours. The same can be seen in the SOFIA data maps from Hankins et al. (2019, 2020). We therefore conclude that it is unlikely that this dust ridge is solely caused by heating from young, massive star-forming regions in the CMZ.
@@ 5.5.1 CMZ Column Densities and Dust Temperatures
We compare our dust temperature and column density measurements with others in the literature, many of which are based on the same dataset but use a variety of techniques. We find overall very good agreement between the absolute values of both dust temperature and column density as well as their overall spatial distribution in comparison with Marsh et al. (2017) and Tang et al. (2021a, b), but find a discrepancy between our column densities and those from Molinari et al. (2011c).
Each of these measurements have advantages and disadvantages. Our work uniquely and uniformly provides the column density and dust temperature measurements across the entire inner 40 longitude of the Galaxy. Because of this, we do not include higher resolution data available only in the CMZ and perform a simple single temperature modified blackbody fit. One advantage of this approach is a comprehensive fore/background subtraction process, applied automatically over the entire inner 40. This fore/background subtraction process was important for our purposes since the following papers in this series are focused on the hierarchical cloud-scale and CMZ-wide properties of the dense gas. However, we enthusiastically refer the reader to other works (particularly Tang et al., 2021a, b; Marsh et al., 2017) for higher resolution column density and temperature maps in just the CMZ.
@@
Our dust temperatures show overall good agreement with Molinari et al. (2011c), Marsh et al. (2017), and Tang et al. (2021a, b) in both the absolute values of the dust temperature across the region and its spatial distribution, including of the warm, dust ridge (Section 5.4). The agreement with Molinari et al. (2011c) and Tang et al. (2021a, b) is especially good in regions of high column density, where the measurements are well within the uncertainties (variation of about 1-2 K). In low column density regions, the variation can be higher (about 4-7 K), with the Tang et al. (2021a, b) temperatures being systematically higher overall, with the exception of a few low column density ‘hot spots’ identified in our maps, which are not found in these maps. This is likely due to variations in the methodology, with this work focusing on careful fore/background subtraction and the Tang et al. (2021a, b) work incorporating new AzTEC data as well as careful modeling to achieve higher spatial resolution. The overall trends in temperature are consistent amongst our maps and the PPMAP results. However, the PPMAP technique finds a uniformly lower dust temperature (about $2-3$ K in high column density regions and up to 10 K in lower column density regions). While this is likely partly due to the higher spatial resolution in the PPMAP images, this cannot be the full explanation, since we would expect not only deeper troughs but also higher peaks if this were purely a resolution effect.
@@
Similarly, the column densities derived show overall good agreement with previous works in the literature (Marsh et al., 2017; Tang et al., 2021a, b). The column densities we report, however, appear to be significantly lower (by up to a factor of ten) than those from Molinari et al. (2011c). Some difference is expected since that work did not subtract a fore/background which can account for up to a factor of two and they also used a relationship between $\tau_{dust}$ and $N_{H}$ that is calibrated at high galactic latitudes which could be a factor of 2-4 higher in the plane (Molinari et al., 2011c). Ponti et al. (2015) also find that column densities derived by Molinari et al. (2011) are larger than their constraints from X-ray modeling by a factor of 5-10. Our column densities agree very well with the work from Tang et al. (2021a, b), with a variation of about 20% in both directions. However, our values are systematically lower than those from Marsh et al. (2017) by about a factor of 2.5. This is likely due to a combination of factors. For one, our resolution is lower, so high column density regions would be smeared out in our maps and have a higher intrinsic column density in the Marsh et al. (2017) approach. Secondly, our careful fore/background subtraction technique removes some contaminating emission which leads to a lower overall value of the column density. Finally, the PPMAP temperatures are lower, which would lead to a higher measured column density for similar flux values. Additionally, each paper used slightly different dust opacities, which can explain some amount of discrepancy in the resulting column densities. As described in more detail in Section 3.1.4, these discrepancies are in line with the intrinsic uncertainty expected in these measurements of about a factor of two in the column densities.
@@ Appendix A Appendix: Data Release
All data analysis products from this work are freely available on the Harvard Dataverse here: https://dataverse.harvard.edu/dataverse/3D_CMZ. We describe here the data products, and this information is also available in the Dataverse metadata.
@@
  • the full dendrogram decomposition (higalcmz_dendrogram.fits) as well as a mask of all the dendrogram structures labeled by their structure IDs (higalcmz_column_dendrogram_mask.fits).

### v2

@@ 5.3 The Total Dense Gas Mass of the CMZ
The column density map of the CMZ can be integrated to estimate the total dense molecular gas mass of the CMZ. As described in Section 3, we use the fore/background subtracted maps for our main analysis, but include constraints from the maps with the fore/background included in Table 1 as well. We consider any gas with a measured column density of N(H2) $>10^{22}$ cm-2 to be dense. In the translation of high column density dust seen with Herschel to the inference of dense, molecular gas are several assumptions: 1) high column density gas corresponds to high volume density gas and 2) dust seen with Herschel traces dense, molecular gas. These assumptions are supported by the many studies (e.g. Mills et al., 2018b; Krieger et al., 2017; Ginsburg et al., 2016) which find abundant dense, molecular gas in the CMZ, with intensity peaks that largely correspond with the Herschel data (Mills & Battersby, 2017). We estimate that most of the gas traced with Herschel observations is likely $n>10^{4}\>\mathrm{cm}^{-3}$ and we refer to this as ‘dense, molecular gas’ throughout.
@@ 5.4 A Ridge of Warm Dust
In the second hypothesis, the ridge of warm dust is thought to be due to heating from young, still-forming or just formed massive stars that follow the morphology of this structure. Detailed maps, including some with temperature, from Hankins et al. (2019, 2020) as well as from the GLIMPSE survey (Benjamin et al., 2003) show mid-IR emission from many young, massive star-forming regions in the CMZ. While there are quite a few located in the central region of the warm dust ridge just above SgrA*, otherwise there seems to be no special association of the young, forming massive star population with the morphology of the dust ridge or other warm spots. In Figure 5 we show the warm dust contours on top of the Spitzer 8 $\mu$m data from the GLIMPSE survey (Benjamin et al., 2003). The central bright massive star-forming regions are clearly seen at 8 $\mu$m, but otherwise, a visual inspection shows no trend with the morphology of the warm dust contours. The same can be seen in the SOFIA data maps from Hankins et al. (2019, 2020). We therefore conclude that it is unlikely that this dust ridge is solely caused by heating from young, massive star-forming regions in the CMZ.
@@ 5.5 Comparison with Previous Work
Figure 7: A quantitative comparison of column densities and temperatures from the literature with this work shows a decent amount of scatter, but an overall good agreement, with the notable exception of the Molinari et al. (2011c) column densities. The left plot shows the column densities computed in this work on the x-axis versus the column densities in Marsh et al. (2017) in dark blue, Tang et al. (2021a, b) in purple, Molinari et al. (2011c) in orange, and the column densities without background subtraction in yellow. Results of linear fits are shown in the dashed lines and expression of the same color. The right plot shows the comparison between the dust temperatures for all regions with N(H2) $>10^{23}$ cm-2 from this work on the x-axis, and other regions in colors that match the left plot. Linear fits are again displayed as dashed lines and expressions of the same color.
@@ 5.5.1 CMZ Column Densities and Dust Temperatures
We compare our dust temperature and column density measurements with others in the literature, Marsh et al. (2017), Tang et al. (2021a, b), Molinari et al. (2011c), which are based on the same dataset but use a variety of techniques. We present a quantitative comparison in Figure 7. The left panel shows the comparison between the column densities, N(H2), along with linear fits and the right panel shows the comparison with dust temperatures along with linear fits. These plots show a decent amount of scatter, but overall good agreement between both dust temperatures and column densities. The notable exception is a large disagreement with the Molinari et al. (2011c) column densities.
To create Figure 7 we retrieved published datasets from the hyperlinks indicated in publications for Marsh et al. (2017) and Tang et al. (2021a, b) and via private communication for Molinari et al. (2011c). We used the package astropy reproject to reproject all of the datasets onto the smallest footprint, which was Tang et al. (2021a, b). We convert all of the data to the same units, remove NaNs, and then plot the pixel-by-pixel comparison between each dataset and the measurements in this work. Since the pixel size is smaller than the beam, these points are not all independent. We use the package numpy polyfit to perform simple linear fits of each dataset relative to the measurements in this work. For both measurements, we only fit a line to points with a column density N(H2) $>10^{22}$ cm-2 since this is approximately the level of the background, making data below this point less useful in comparison.
Each of these measurements have advantages and disadvantages. Our work uniquely and uniformly provides the column density and dust temperature measurements across the entire inner 40 longitude of the Galaxy, therefore we do not include higher resolution data available only in the CMZ. One advantage of this approach is a comprehensive fore/background subtraction process, applied automatically over the entire inner 40. This fore/background subtraction process was important for our purposes since the following papers in this series are focused on the hierarchical cloud-scale and CMZ-wide properties of the dense gas. However, we enthusiastically refer the reader to other works (particularly Tang et al., 2021a, b; Marsh et al., 2017) for higher resolution column density and temperature maps in just the CMZ.
@@
As seen in the right panel of Figure 7, dust temperatures show overall good agreement, including of the warm, dust ridge (Section 5.4). In addition to the linear fits shown on the plot, we find that the overall ratio between each literature temperatures and our temperatures (e.g. Tang temp / Battersby temp = ratio) is as follows: the Tang et al. (2021a, b) ratio has a median value of 1.07 and a standard deviation of 0.08, the Molinari et al. (2011c) ratio is 1.06 $\pm$ 0.03, the with background temperatures have a ratio of 0.98 $\pm$ 0.03, and the Marsh et al. (2017) ratio is 0.87 $\pm$ 0.04. We note that the agreement with Molinari et al. (2011c) and Tang et al. (2021a, b) is especially good in regions of high column density, but in low column density regions, the variation can be higher, with the Tang et al. (2021a, b) temperatures being systematically higher overall. The overall trends in temperature are consistent among our maps and the PPMAP results. However, the PPMAP technique finds a uniformly lower dust temperature. While this is likely partly due to the higher spatial resolution in the PPMAP images, this cannot be the full explanation, since we would expect not only deeper troughs but also higher peaks if this were purely a resolution effect.
@@
As seen in the left panel of Figure 7, the column densities derived show overall reasonable agreement with previous works in the literature (Marsh et al., 2017; Tang et al., 2021a, b). In addition to the linear fits shown on the plot, we find that the overall ratio between each literature column density and our column density (e.g. Tang column density / Battersby column density = ratio) is as follows: the Tang et al. (2021a, b) ratio has a median value of 0.84 and a standard deviation of 0.18, the Molinari et al. (2011c) ratio is 11.4 $\pm$ 1.4, the with background column densities have a ratio of 1.41 $\pm$ 0.21, and the Marsh et al. (2017) ratio is 2.1 $\pm$ 0.43. The column densities we report, are significantly lower than those from Molinari et al. (2011c). Some difference is expected since that work did not subtract a fore/background which can account for up to a factor of two and they also used a relationship between $\tau_{dust}$ and $N_{H}$ that is calibrated at high galactic latitudes which could be a factor of 2-4 higher in the plane (Molinari et al., 2011c). Ponti et al. (2015) also find that column densities derived by Molinari et al. (2011) are larger than their constraints from X-ray modeling by a factor of 5-10. Our values agree with reasonable certainty with the Tang et al. (2021a, b) but our values are systematically lower than those from Marsh et al. (2017) by about a factor of 2. This is likely due to a combination of factors. For one, our resolution is lower, so high column density regions would be smeared out in our maps and have a higher intrinsic column density in the Marsh et al. (2017) approach. Secondly, our careful fore/background subtraction technique removes some contaminating emission which leads to a lower overall value of the column density. Finally, the PPMAP temperatures are lower, which would lead to a higher measured column density for similar flux values. Additionally, each paper used slightly different dust opacities, which can explain some amount of discrepancy in the resulting column densities. As described in more detail in Section 3.1.4, these discrepancies (with the exception of the Molinari et al. (2011c) column densities) are in line with the intrinsic uncertainty expected in these measurements of about a factor of two in the column densities.
@@ Appendix A Appendix: Data Release
All data analysis products from this work are freely available on the Harvard Dataverse here: https://dataverse.harvard.edu/dataverse/3D_CMZ. The dataset for this paper is found in the Papers I and II repository: https://doi.org/10.7910/DVN/7DOJG5 (Battersby et al., 2024). We describe here the data products, and this information is also available in the Dataverse metadata.